### v1

@@ Abstract
Lung cancer remains one of the leading causes of morbidity and mortality worldwide, making early diagnosis critical for improving therapeutic outcomes and patient prognosis. Computer-aided diagnosis (CAD) systems, which analyze CT images, have proven effective in detecting and classifying pulmonary nodules, significantly enhancing the detection rate of early-stage lung cancer. Although traditional machine learning algorithms have been valuable, they exhibit limitations in handling complex sample data. The recent emergence of deep learning has revolutionized medical image analysis, driving substantial advancements in this field. This review focuses on recent progress in deep learning for pulmonary nodule detection, segmentation, and classification. Traditional machine learning methods, such as SVM and KNN, have shown limitations, paving the way for advanced approaches like Convolutional Neural Networks (CNN), Recurrent Neural Networks (RNN), and Generative Adversarial Networks (GAN). The integration of ensemble models and novel techniques is also discussed, emphasizing the latest developments in lung cancer diagnosis. Deep learning algorithms, combined with various analytical techniques, have markedly improved the accuracy and efficiency of pulmonary nodule analysis, surpassing traditional methods, particularly in nodule classification. Although challenges remain, continuous technological advancements are expected to further strengthen the role of deep learning in medical diagnostics, especially for early lung cancer detection and diagnosis. A comprehensive list of lung cancer detection models reviewed in this work is available at https://github.com/CaiGuoHui123/Awesome-Lung-Cancer-Detection.
@@ 1 Introduction
Lung cancer is one of the leading causes of both incidence and mortality worldwide. According to the latest assessment by the International Agency for Research on Cancer (IARC) of the World Health Organization (WHO), in 2022, there were approximately 19.96 million new cancer cases globally, with 2.48 million of them being lung cancer, accounting for 12.4% of all cases, making it the most prevalent cancer worldwide Fig. 1 [3]. Due to the subtlety of early-stage lung cancer symptoms, patients often miss the optimal treatment window. Early screening and CT scans are essential for early diagnosis, as they allow for the detection of lesions before clinical symptoms manifest, significantly increasing the rate of early lung cancer detection. Computer-aided diagnosis (CAD) systems are designed to assist in medical image analysis by automating the detection of abnormalities, such as those in CT scans [4]. These systems improve diagnostic accuracy and reduce the workload of radiologists. In lung cancer detection, CAD systems are particularly effective in identifying and assessing pulmonary nodules. With advances in deep learning and artificial intelligence, these systems have become increasingly accurate, fast, and reliable, supporting early lung cancer diagnosis.
Fig. 1: Pie chart showing the distribution of cases and deaths for the top five cancers for both sexes in 2022. The area of the pie chart reflects the proportion of the total number of cases or deaths [3].
@@ 2.4 Generative Adversarial Network (GAN)
Moreover, Tyagi et al. [53] proposed a 3D Conditional GAN (CSE-GAN) for pulmonary nodule segmentation, integrating a U-Net-based generator with parallel compression and excitation modules. This approach effectively reduced overfitting and improved segmentation performance, achieving a Dice coefficient of 80.74% on the LUNA16 dataset. These advanced GAN-based methods not only enhance training datasets with high-quality synthetic images but also improve segmentation and classification accuracy, effectively addressing issues related to data scarcity and class imbalance.
@@ 5.1 Discussion on AI-Driven Advances in Lung Cancer Detection
Key Innovations and Performance Comparisons: Over the past several years, significant advancements have been made in AI-driven pulmonary nodule detection, contributing uniquely to the evolution of the field. In 2020, Masud et al. [90] developed a lightweight deep learning model achieving 97.9% accuracy, showcasing the potential for AI deployment in resource-constrained environments, such as mobile devices in clinics with limited computational power. In 2022, Bhatt et al. [22] demonstrated the utility of YOLOv4 for real-time detection with 95% precision, emphasizing the practical application of AI in rapid diagnostics. Fu et al. [37] integrated 3D CT imaging with serum biomarkers, highlighting a multimodal approach that combines diverse data sources to enhance diagnostic precision. In 2023, Gugulothu et al. [51] introduced a hybrid deep learning approach, and by 2024, Gautam et al. [88] combined ResNet, DenseNet, and EfficientNet to achieve a sensitivity of 98.6%, significantly reducing false negatives. Furthermore, Usman et al. [70] developed MEDS-Net, which achieved a CPM score of 93.6% by reducing false positives, thus minimizing unnecessary invasive procedures and enhancing diagnostic reliability. These innovations collectively advance the clinical applicability of AI for lung cancer detection by improving both accuracy and efficiency.

### v2

@@ Abstract
Lung cancer remains one of the leading causes of morbidity and mortality worldwide, making early diagnosis critical for improving therapeutic outcomes and patient prognosis. Computer-aided diagnosis systems, which analyze computed tomography images, have proven effective in detecting and classifying pulmonary nodules, significantly enhancing the detection rate of early-stage lung cancer. Although traditional machine learning algorithms have been valuable, they exhibit limitations in handling complex sample data. The recent emergence of deep learning has revolutionized medical image analysis, driving substantial advancements in this field. This review focuses on recent progress in deep learning for pulmonary nodule detection, segmentation, and classification. Traditional machine learning methods, such as support vector machines and k-nearest neighbors, have shown limitations, paving the way for advanced approaches like Convolutional Neural Networks, Recurrent Neural Networks, and Generative Adversarial Networks. The integration of ensemble models and novel techniques is also discussed, emphasizing the latest developments in lung cancer diagnosis. Deep learning algorithms, combined with various analytical techniques, have markedly improved the accuracy and efficiency of pulmonary nodule analysis, surpassing traditional methods, particularly in nodule classification. Although challenges remain, continuous technological advancements are expected to further strengthen the role of deep learning in medical diagnostics, especially for early lung cancer detection and diagnosis. A comprehensive list of lung cancer detection models reviewed in this work is available at https://github.com/CaiGuoHui123/Awesome-Lung-Cancer-Detection.
@@ 1 Introduction
Lung cancer is one of the leading causes of both incidence and mortality worldwide. According to the latest assessment by the International Agency for Research on Cancer (IARC) of the World Health Organization (WHO), in 2022, there were approximately 19.96 million new cancer cases globally, with 2.48 million of them being lung cancer, accounting for 12.4% of all cases, making it the most prevalent cancer worldwide Fig. 1 [1]. Due to the subtlety of early-stage lung cancer symptoms, patients often miss the optimal treatment window. Early screening and computed tomography (CT) scans are essential for early diagnosis, as they allow for the detection of lesions before clinical symptoms manifest, significantly increasing the rate of early lung cancer detection. Computer-aided diagnosis (CAD) systems are designed to assist in medical image analysis by automating the detection of abnormalities, such as those in CT scans [2]. These systems improve diagnostic accuracy and reduce the workload of radiologists. In lung cancer detection, CAD systems are particularly effective in identifying and assessing pulmonary nodules. With advances in deep learning and artificial intelligence, these systems have become increasingly accurate, fast, and reliable, supporting early lung cancer diagnosis.
Fig. 1: Pie chart showing the distribution of cases and deaths in 2022 for the top five cancers and an aggregated “Others” category (all remaining cancer types). The area of each slice reflects its proportion of the total number of cases or deaths [1].
@@ 2.4 Generative Adversarial Network (GAN)
Moreover, Tyagi et al. [53] proposed a 3D Conditional GAN (CSE-GAN) for pulmonary nodule segmentation, integrating a U-Net-based generator with parallel compression and excitation modules. This approach effectively reduced overfitting and improved segmentation performance, achieving a Dice coefficient of 80.74% on the LUNA16 dataset. These advanced GAN-based methods not only enhance training datasets with high-quality synthetic images but also improve segmentation and classification accuracy, effectively addressing issues related to data scarcity and class imbalance. On the other hand, we have greatly enriched our discussion of Generative Adversarial Networks (GANs) by detailing not only their applications in data augmentation but also the quantitative metrics used to assess the quality and realism of synthetic medical images. Key evaluation measures now include the Fréchet Inception Distance (FID), which quantifies distributional similarity between generated and real images in the feature space of a pretrained network (lower is better); the Inception Score (IS), which captures both image quality and diversity via entropy-based measures (higher is better); the Structural Similarity Index Measure (SSIM), commonly used in medical imaging to assess perceptual similarity between synthetic and original CT slices; and the Peak Signal-to-Noise Ratio (PSNR), which evaluates reconstruction fidelity. We further introduce the concept of a Visual Turing Test (VTT), wherein expert radiologists attempt to distinguish synthetic from real images as a subjective but clinically pertinent validation. Finally, we reference recent studies—such as Gugulothu et al. (2023)—that employ GANs for both image synthesis and enhancement, demonstrating improvements in downstream classification performance when augmented with high-quality synthetic data.
@@ 4 Datasets and Benchmarks
PN9 8798 Nodule Detection & Classification Large-scale Model Training & Validation
@@
The PN9 dataset [83], released in 2020, is a relatively new and valuable resource for pulmonary nodule analysis. It contains high-resolution CT images collected from multiple institutions and scanners, enhancing its diversity and representativeness. One of its key strengths lies in the inclusion of subtle and diagnostically challenging nodules, which are often underrepresented in older datasets. Additionally, PN9 offers clinical annotations that are closely aligned with contemporary diagnostic standards, making it particularly suitable for training and validating models intended for real-world clinical deployment.
@@
We also conducted a detailed analysis of distributional imbalances in the two most widely used public datasets, LIDC-IDRI and LUNA16. In LIDC-IDRI, benign nodules account for 82% (3,572 / 4,356) of all annotated nodules, leading to a strong class skew. Gautam et al. (2024) have shown that models trained on this imbalanced distribution suffer a 12.8% drop in sensitivity when evaluated on clinical cohorts with a more realistic 45% benign-to-malignant ratio. Moreover, both datasets exhibit a size imbalance, with the vast majority of nodules falling in the 3–10 mm range and very few examples of very small or large nodules; a label imbalance, where certain subtypes (e.g., part-solid or ground-glass) are underrepresented; and an anatomical imbalance, as nodules cluster in specific lung regions. To help mitigate some of these issues, we have updated Table  2 to include the more recently released PN9 (2020) dataset, which provides a larger and more evenly distributed sample across nodule sizes, subtypes, and acquisition sources.
@@ 5.1 Discussion on AI-Driven Advances in Lung Cancer Detection
Key Innovations and Performance Comparisons: Over the past several years, significant advancements have been made in AI-driven pulmonary nodule detection, contributing uniquely to the evolution of the field. In 2020, Masud et al. [91] developed a lightweight deep learning model achieving 97.9% accuracy, showcasing the potential for AI deployment in resource-constrained environments, such as mobile devices in clinics with limited computational power. In 2022, Bhatt et al. [22] demonstrated the utility of YOLOv4 for real-time detection with 95% precision, emphasizing the practical application of AI in rapid diagnostics. Fu et al. [37] integrated 3D CT imaging with serum biomarkers, highlighting a multimodal approach that combines diverse data sources to enhance diagnostic precision. In 2023, Gugulothu et al. [51] introduced a hybrid deep learning approach, and by 2024, Gautam et al. [89] combined ResNet, DenseNet, and EfficientNet to achieve a sensitivity of 98.6%, significantly reducing false negatives. Furthermore, Usman et al. [70] developed MEDS-Net, which achieved a CPM score of 93.6% by reducing false positives, thus minimizing unnecessary invasive procedures and enhancing diagnostic reliability. These innovations collectively advance the clinical applicability of AI for lung cancer detection by improving both accuracy and efficiency. To address dataset imbalances, recent studies have adopted dynamic weighted loss functions and carefully tuned oversampling strategies. Dynamic weighting—where training losses for minority classes are up-weighted on the fly—has been shown to improve recall on underrepresented nodules, reducing false negatives without destabilizing overall training. However, overly aggressive oversampling in LUNA16 can introduce synthetic bias and lead to overfitting, ultimately harming model generalization on unseen data. We therefore highlight a balanced approach: combining moderate oversampling with cost-sensitive learning and regularization techniques to narrow the generalization gap and mitigate diagnostic blind spots in real-world clinical deployments.
@@
Single-Stage Detectors for Real-Time Nodule Analysis: Single-stage object detectors such as YOLO-v5 and YOLO-v6 have emerged as powerful tools that balance high detection accuracy with real-time inference, making them ideally suited for integration into clinical workflows. YOLO-v5 bypasses the region-proposal stage by predicting bounding boxes and class probabilities in a single network pass; for instance, Ahmadyar et al. (2024) combined a hierarchical YOLO-v5s with a 3D CNN backbone on LUNA16, achieving 97.8 % sensitivity at a 0.3 confidence threshold [29]. Wu et al. (2024) further refined this framework with their YOLO-MSRF variant—incorporating multi-scale feature refinement—to reach 94.02 % sensitivity and 95.26 % mAP on the same benchmark [97]. YOLO-v6 further advances the architecture by employing an optimized backbone and anchor-free detection head to reduce latency and enhance small-object recall—key for identifying subtle or diminutive nodules—and early reports indicate over 98 % sensitivity on LIDC-IDRI with per-slice inference times under 50 ms. These developments underscore a trend toward embedding lightweight, fast, and accurate CAD tools directly within PACS or imaging consoles, thereby broadening access to AI-driven lung cancer screening in routine clinical practice.
@@ 5.3 Discussion on AI-Driven Advances in Lung Cancer Classification
Subtyping nodules enables more precise clinical decisions: we specifically discuss the subclassification of lung nodules—beyond the traditional benign/malignant dichotomy—to emphasize its clinical significance. Ground-glass nodules (GGNs), for example, are often early indicators of adenocarcinoma, whereas calcified nodules tend to be benign; recognizing these differences can directly influence follow-up intervals, biopsy decisions, and surgical planning. Recent studies demonstrate the feasibility of this finer-grained analysis: Naveen et al. (2023) report subtype-specific performance on LIDC-IDRI of 90.86% for solid, 93.68% for part-solid, and 97.06% for non-solid nodules; Guo et al. (2023) and Sivakumar et al. (2024) achieve over 95% accuracy on benign versus malignant classification, providing a strong foundation for multi-class extensions; and Zhang et al. (2021, 2022) incorporate segmentation and multi-task learning to fuse morphological and contextual features, further improving subtype discrimination. These advancements in multi-class classification and multi-task frameworks offer a more nuanced and clinically actionable approach to pulmonary CAD systems.
FDA clearance validates AI lung CAD for clinical use: we have added the following discussion of FDA-cleared AI systems to illustrate their clinical translation: Several AI systems for lung nodule detection and classification have obtained FDA clearance, demonstrating the readiness of deep learning–based CAD tools for clinical deployment. Optellum Lung Cancer Prediction (2021) leverages a deep learning–derived “lung cancer prediction score” trained on over 300,000 CT scans; a retrospective study in Radiology (2021) reported a 12–20 % increase in pulmonologist diagnostic accuracy and a significant reduction in unnecessary follow-ups. Arterys Lung AI (2020) provides automated nodule detection, classification, and volumetric tracking via cloud-PACS integration, reporting robust performance on multi-institutional datasets for both solid and subsolid nodules. Although CT-based systems dominate, Lunit INSIGHT CXR (2021) earned FDA clearance for chest X-ray–based detection of ten thoracic abnormalities—including nodules—achieving an AUC of 0.974 in a clinical validation of over 100,000 images. Finally, VUNO Med-LungCT AI, cleared under Korea’s 510(k) pathway and pending FDA approval, has demonstrated > 90 % sensitivity in multicenter Korean CT cohorts and is now integrated into early lung-cancer screening workflows. These examples underscore the importance of explainability, regulatory compliance, and seamless integration into clinical systems when translating deep learning methods from research to routine practice.